Q1 Valentino’s Stage Productions is taking on a new project. The project is expected to increase net income by $1,000,000 for each of the next 3 years. The equipment needed will cost $7,000,000. Valentino’s hurdle rate is 12% and has the following capital structure.

Bond issuance:

20% of total funds, requires 15% interest per year

Bank loan:

60% of total funds, requires 9.5% interest per year

Preferred Stock issuance:

20% of total funds, requires 5% dividend per year

What is Valentino’s average rate of return on the new project?

a. 9.29%

b. 11.29%

c. 28.57%

d. 40.00%

Q2 Valentino’s Stage Productions is taking on a new project. The project is expected to increase net income by $1,000,000 for each of the next 3 years. The equipment needed will cost $7,000,000. Valentino’s hurdle rate is 12% and has the following capital structure.

Bond issuance:

20% of total funds, requires 15% interest per year

Bank loan:

60% of total funds, requires 9.5% interest per year

Preferred Stock issuance:

20% of total funds, requires 5% dividend per year

Should Valentino invest in the project?

a. No, the average rate of return does not exceed the hurdle rate

b. No, the average rate of return exceeds the hurdle rate

c. Yes, the average rate of return does not exceed the hurdle rate

d. Yes, the average rate of return exceeds the hurdle rate

Q3 Robinson Bottling Co. reported Sales of $220,000, Cost of Goods Sold of $87,000 (including depreciation expense of $10,000), and Office Expenses of $11,000. The tax rate is 15%. The company has no significant changes in Accounts Receivable, Inventory, Accounts Payable, Prepaid Expenses, or Unearned Revenue.

What are Robinson’s net after-tax cash flows?

a. $103,700

b. $105,200

c. $110,800

d. $113,700
Answer #1
Q1] Average rate of return = Annual net income/Average investment = 1000000/(7000000/2) = 28.57%
Answer: [c] 28.57%
Q2] Should Valentino invest in the project?
Answer: [d] Yes, the average rate of return exceeds the hurdle rate
Q3] Net income = (220000-87000-11000)*(1-15%) = $          103,700
Net after tax cash flows = NI+Depreciation = 103700+10000 = $          113,700
Answer: [d] $113,700